Circular No. 8219

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

POSSIBLE NOVA IN CIRCINUS W. Liller, Viña del Mar, Chile, reports his discovery of a possible nova

on Technical Pan film images taken with a 85-mm Nikon camera lens (+ red filter) on Oct. 9.004 (mag ≈ 7.7) and 9.008 UT (mag ≈ 8.2). The new object, which was absent on exposures taken on Sept. 26.004 (limiting mag 11.5) and Oct. 4.010 (limiting mag 11.0), is located at α = 15^{h}17^{m}.8, δ = −61^{o}58′ (equinox 2000.0).

SUPERNOVAE 2003hl AND 2003iq IN NGC 772 Jean-Marie Llapasset, Perpignan, France, reports the discovery of an

apparent supernova (mag 16.4), on 26 unfiltered CCD images taken with a 0.28-m reflector around Oct. 8.958 UT, while monitoring SN 2003hl (cf. IAUC 8184), which itself was at mag ∼ 16.5. The new object is located at α = 1^{h}59^{m}19^{s}.96, δ = +18^{o}59′42′′.1 (equinox 2000.0), which is 5′′.3 east and 45′′.9 south of the nucleus of NGC 772, and ∼ 37′′ southwest of SN 2003hl. H. Yamaoka, Kyushu University, reports that R-band images taken by K. Ayani, Bisei Astronomical Observatory (BAO), on Oct. 9.60 with BAO 1.01-m telescope show the new object at R ∼ 16.0 and at position end figures 19^{s}.92, 41′′.8. An image taken by Llapasset on Oct. 5.97 does not show SN 2003iq (limiting mag ∼ 17.5). Additional magnitudes of SN 2003hl by Llapasset: Aug. 28.182, 16.6, Oct. 5.965, 16.8.

SUPERNOVA 2000gb Further to IAUC 8218, D. S. Madgwick adds that another type-Ia

supernova has been detected in the first SDSS data release, located at α = 17^{h}32^{m}28^{s}.54, δ = +56^{o}04′25′′.5 (equinox 2000.0), with a redshift of 0.1225 and an estimated magnitude of V = 20.4 on 2000 Oct. 1.

AG DRACONIS E. O. Waagen, AAVSO, reports that the current outburst of AG Dra

(cf. IAUC 8217) is another in a series of smaller outbursts that have occurred since that last major outburst in 1991, the most recent of them occurring in Sept. 2002 (when the star reached m_{v} ∼ 9.2). AAVSO visual magnitude estimates: Sept. 25.868 UT, 10.1 (E. Broens, Mol, Belgium); 28.853, 9.7 (P. Cloesen, Beringen, Belgium); 29.795, 9.2 (E. Muyllaert, Oostende, Belgium); Oct. 1.158, 8.8 (R. Royer, Springville, CA); 3.970, 9.4 (B. Granslo, Fjellhamar, Norway); 8.000, 9.2 (Granslo).

2003 October 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green